# Is animal magnetoreception quantum? A perspective from the energy resolution limit

### v1

@@ VI Global ER MAP for biological magnetoreception
In Fig. 6 we present the global map of the spectral magnetic sensitivity given by $\widetilde{\delta B}=\delta B\sqrt{\tau}$ as a function of the sensor’s linear dimension $L$ for all sensing mechanisms we have considered. We remind the reader that we used a common value for the magnetic sensitivity, $\delta B=0.3~{}\upmu{\rm T}$. The parameters that differ for each sensing mechanism are the measurement time $\tau$ and the linear dimension $L$. From the global map we can draw a number of conclusions.
From the perspective of the ERL, the radical-pair magnetoreception can indeed be “quantum”, if the number $N_{\rm cry}$ of cryptochromes participating in the sensing is $N_{\rm cry}=N_{\rm cry}^{0}=2\times 10^{5}$ (which number depends on the specific values of $\delta B$ and $\tau$). For $N_{\rm cry}>N_{\rm cry}^{0}$, the energy resolution grows beyond $\hbar$, and the “quantumness” is reduced. Similar comments apply to the MagR sensor, with the relevant minimum number of contributing complexes being $N_{\rm MagR}^{0}=12$.
The magnetite mechanism, and even more so the CPF measurement appear to be more “classical”, which is probably not surprising. Nevertheless, the ER of the magnetite mechanism stands at the level of $10^{6}\hbar$, which is not too far from the quantum limit. As interactions of magnetized materials with the magnetic field are at their core quantum, one could imagine pushing the magnetite mechanism closer to the quantum limit with biomimetic engineering.
@@ VII Conclusions
In particular, the most obvious utility of the energy resolution limit is to extract information about unknown system parameters. For example, given the magnetic sensitivity and the measurement time of the cryptochrome and MagR magnetoreceptors, we extract for the first time the minimum number of magnetoreceptors required for sensing, since it is this minimum number that is required to satisfy the ERL bound. Conversely, the fact that a given magnetic sensing model might satisfy the ERL, as for example the magnetite model does, is not proof that the model is correct, but merely shows it is a physically viable model.

### v2

@@ VI Global ER MAP for biological magnetoreception
In Fig. 6 we present the global map of the spectral magnetic sensitivity given by $\widetilde{\delta B}=\delta B\sqrt{\tau}$ as a function of the sensor’s linear dimension $L$ for all three sensing mechanisms we have considered. We remind the reader that we used a common value for the magnetic sensitivity, $\delta B=0.3~{}\upmu{\rm T}$. The parameters that differ for each sensing mechanism are the measurement time $\tau$ and the linear dimension $L$. From the global map we can draw a number of conclusions.
From the perspective of the ERL, the radical-pair magnetoreception can indeed be “quantum”, if the number $N_{\rm cry}$ of Cry proteins participating in the sensing equals $N_{\rm cry}^{0}(\tilde{\tau})=2\times 10^{5}/\tilde{\tau}$, which number depends on the specific value of $\delta B$ in vivo, the parameter $\tilde{\tau}$ quantifying spin relaxation time, and the sensor’s volume. For $N_{\rm cry}>N_{\rm cry}^{0}(\tilde{\tau})$, the energy resolution grows beyond $\hbar$, and the “quantumness” is reduced. Similar comments apply to the MagR sensor, with the relevant minimum number of contributing complexes being $N_{\rm MagR}^{0}(\tilde{\tau})=10/\tilde{\tau}$.
The magnetite mechanism, and even more so the CPF measurement appear to be more “classical”, which is probably not surprising. Nevertheless, the ER of the magnetite mechanism stands at the level of $10^{6}\hbar$, which is not too far from the quantum limit. As interactions of magnetized materials with the magnetic field are at their core quantum, the prospect of pushing the magnetite mechanism closer to the quantum limit with biomimetic engineering seems reasonable.
@@ VII Conclusions
In particular, the most obvious utility of the energy resolution limit is to extract information about unknown system parameters. For example, given the magnetic sensitivity and the measurement time of the Cry and MagR magnetoreceptors, we extract for the first time the minimum number of magnetoreceptors required for sensing, since it is this minimum number that is required to satisfy the ERL bound. Conversely, the fact that a given magnetic sensing model might satisfy the ERL, as for example the magnetite model does, is not proof that the model is correct, but merely shows it is a physically viable model.